# An electron-hadron collider at the high-luminosity LHC

### v1

@@ 2 The LHeC interaction region
It should be emphasised that the case described here refers to the truly concurrent operation of electron-hadron and hadron-hadron collisions. So, the aperture needed in the mini-beta quadrupoles of the protons is defined by both the colliding and the non-colliding proton beam. Therefore the “other”, non-colliding proton beam still has to be kept in the storage ring and guided through the $ep$ interaction region untouched. However, in order to save free aperture for the optics of the colliding protons, relaxed beam optics has been developed which allows for minimum beam size and aperture need in the mini-beta quadrupoles: $\beta^{\ast}=24~{}{\rm m}$ has been found as optimum value for the non-colliding proton in this case.
The same collision scheme will be applied for the other electron-hadron collisions. For example, assuming the nominal lead beam parameters at the HL-LHC, i.e. the beam radius of 17 $\mu$m and the $Pb$ bunch population of $1.8\cdot 10^{8}$, one expects the $ePb$ luminosity at the LHeC of $2\cdot 10^{30}~{}{\rm cm}^{-2}{\rm s}^{-1}$, or about $4\cdot 10^{32}~{}{\rm cm}^{-2}{\rm s}^{-1}$ per nucleon.
Polarised electron beams can rather naturally incorporated in an ERL type accelerator. Laser driven state of the art electron guns provide polarised beams up to 90% as previously suggested in the LHeC CDR [1]). Harmful effects like resonant depolarisation that play a major role in storage rings are by definition not existent and the de-tuning of the spin precession due to the emitted synchrotron light is estimated to be on the order of a few percent. As a consequence the acceleration chain provided by the ERL is quasi spin transparent and does not play a large role in the final – longitudinal – electron beam polarisation at the IP.
@@ 3 Physics case
In phase-one LHeC, the centre-of-mass energy for electron-proton collisions $\sqrt{s_{ep}}=\sqrt{4E_{e}E_{p}}=0.75$ TeV, in place of 1.2 TeV for the 50 GeV electron beam at the final LHeC. This means that the kinematic reach on the $x-Q^{2}$ plane will already be greatly extended with respect to the data taken at HERA at $\sqrt{s}=0.3$ TeV, as $Q^{2}_{\rm max}\simeq s_{ep}$ and $x_{\rm min}\propto 1/s_{ep}$. As discussed in the LHeC CDR [1], a relatively low integrated luminosity of $50~{}{\rm fb}^{-1}$ is sufficient to greatly improve the constraints on the parton density functions, particularly at low $x$. Therefore, already after two years of running the phase-one LHeC it will be possible to provide valuable feedback to hadron-hadron experiments at the HL-LHC, what will result in significant improvements in a number of precision measurements [7]. Moreover, with entering into a new kinematic region of very low $x$ the novel QCD phenomena are expected to be observed – new parton dynamics and non-linear parton density effects.
Table 3: Inclusive Higgs, single top, and top pair generator level production cross sections for the phase-one and final LHeC, and for un-polarised and polarised electron beams. All cross sections are quoted at leading order, with the Higgs boson mass of $125~{}{\rm GeV}$ and the top quark mass of $172~{}{\rm GeV}$.
@@
To illustrate the scientific potential of the research in electroweak and Higgs physics at the phase-one LHeC, a few MadGraph5_aMC@NLO [8] generator-level cross sections for un-polarised and longitudinally polarised electron beams are compared in Tab. 3 to those for the final LHeC. The Higgs boson case remains compelling at 0.75 TeV (especially for the charged current events), as well as the single-top-quark one, not only for mastering pioneering electron-proton measurement techniques but also for making measurements of some specific channels. For example, assuming the Run5 total integrated luminosity of $250~{}{\rm fb}^{-1}$ at 0.75 TeV, and an 80% electron longitudinal polarisation, one obtains the following uncertainties by scaling up the errors obtained for the 1.2 TeV LHeC [1] using the numbers of produced events: $\delta\kappa_{W}\approx 2.5\%$, for the Higgs boson coupling to $W$ bosons, and $\delta{\cal R}e(f^{2}_{L})\approx 0.2$, for the anomalous $Wtb$ coupling – both errors comparable to the expected final uncertainties for the corresponding measurements at the HL-LHC using $3~{}{\rm ab}^{-1}$ of the proton-proton data.
@@
In Fig. 6 we show the differential cross-sections in rapidity for the Higgs boson and a single top quark produced at the phase-one LHeC as well as the final one – for the latter the cross-sections peak at about $Y=2$ requiring a good coverage in the forward part of the hadron hemisphere.
@@
Figure 6: Differential cross-sections in rapidity for the Higgs boson (left) and a single top quark (right), both produced through charged current, assuming the phase-one electron energy of 20 GeV, with and without electron polarisation, as well as for the final energy of 50 GeV.
@@
Experimental conditions at the LHeC will be particularly favourable for studies of the exclusive production via photon-photon fusion, the event pile-up will be negligible, highly improving the efficiency and purity of exclusive event selection, and at the same time some sources of overwhelming backgrounds, as due to the Drell-Yan process for example, will be absent. Moreover, no event suppression will occur due to hadronic re-scattering and the associated significant theoretical and experimental uncertainties will not intervene. Finally, data streaming (or the absence of triggering) will allow for detection and reconstruction of a much wider range of final states. As a result, despite the lower average $\gamma\gamma$ centre-of-mass energy than that at the HL-LHC, a significantly larger statistics of detected photon-photon interactions is expected already at the phase-one LHeC.
The prime example is the two-photon production of $W$-boson pairs: assuming an integrated $ep$ luminosity of $250~{}{\rm fb}^{-1}$, about 5000 pairs will be produced at 0.75 TeV [9]. As the semi-leptonic decays will be detected with high efficiency, that will lead to much improved sensitivities to anomalous couplings with respect to the measurements at the HL-LHC.
@@
A fully exclusive two-photon production of tau-lepton pairs has a large cross section of about 40 pb at the phase-one LHeC, for pair invariant masses above 10 GeV [9]. This will lead to very large event statistics within the acceptance of the central detector and as a consequence to excellent sensitivity to the $\tau$ anomalous magnetic moment $a_{\tau}$. With an integrated $ep$ luminosity of 250 fb-1, the expected sensitivity is an order of magnitude better than that achieved at the phase-one LHC. As a result, for the first time the experimental uncertainties will be better than the higher-order corrections to the $\tau$ magnetic moment in the Standard Model. Similarly, large improvements will be achieved in constraining the electric dipole moment of tau-leptons.
@@
Last but not least, already at the phase-one LHeC, also one of the key questions in nuclear physics can be addressed: how the nucleon structure is modified when immersed in a nuclear medium. For example, by a precise measurement and complete unfolding of the distributions of all parton species in a single nucleus comparable to that in the proton, the LHeC will revolutionise our understanding of nuclear structure in crucial and unique kinematic regions [1, 7]. In addition, it will allow to clarify the dynamics for particle production in nuclear collisions in the region of relevance for the LHC. Finally, it should be stressed that the phase-one LHeC provides an important opportunity to perform $eA$ and $AA$ measurements using the same detector.
@@ 4 General detector requirements
In 2022, an extension of the ALICE experiment beyond Run4 at the HL-LHC was proposed in a letter of intent [10]. In that document, two configurations of magnetic field were considered for the new ALICE 3 detector at the IP2 – produced either by the cost effective 7.5 m long solenoid, or by the high-performance combination of a 2.5 m solenoid at the centre with two 2 m long dipoles on its each side. As discussed above, to allow for electron-hadron collisions, the dipole field is necessary for the separation of the electron beam, so the latter approach is desirable in this context.
The most recent ALICE 3 design includes a good-performance tracking detector that covers a wide rapidity range of $|\eta|<4$ [11], which is well compatible with the needs for electron-hadron measurements at the stage-one LHeC. However, in contrast, the present design lacks a hadronic calorimeter (HCAL) while a performant one is obligatory for achieving proper measurements of the DIS charged current events and reconstruction of the decays of top quark, Higgs, $W$ and $Z$ bosons. Moreover, an extension of the electromagnetic calorimeter is desirable in the electron forward direction to ensure a good reconstruction of DIS neutral current events, and in addition, important studies of exclusive production at the LHeC will largely benefit from specialised near-beam detectors in very forward electron and proton directions.
Finally, it should be stressed that the ALICE 3 design already contains a full data streaming mechanism (i.e. a ”trigger-less” data acquisition), which is also very profitable for measurements at the LHeC.
In summary, it seems that the proposed ALICE 3 detector requires relatively small modifications, apart from the addition of HCAL, to make it also appropriate for the phase-one LHeC experiment – in this way, allowing to vastly increase the scientific scope and impact of research performed using the Run5 data.
@@ 5 Conclusions
In this paper, we argue that the phased implementation of LHeC is feasible and offers the best approach. Firstly, in this case the project financial load will be distributed over more years. Secondly, the experience acquired during the first stage of the project will allow a better design of the final LHeC and possibly also allow for its earlier startup. Finally, phase-one LHeC will ensure very important and timely scientific feedback to the hadron-hadron experiments at the HL-LHC, already during Run5.
The conceptual studies presented above demonstrated the high luminosity performance of such a phase-one LHeC. This in turn guarantees excellent scientific output, including very interesting comparative studies of the $eA$ and $AA$ interactions, as the proposed IR design allows for both types of collisions at IP2.
The new detector at the P2 site can be designed by extension of the ALICE 3 proposal, which offers strong synergies and benefits in several aspects – for example, it is expected to result in its better performance and richer scientific output for studies of $AA$ interactions.
@@ Acknowledgements
L. Forthomme and K. Piotrzkowski appreciate the financial support of the Polish National Agency for Academic Exchange (NAWA) under grant number BPN/PPO/2021/1/00011. L. Forthomme has also been supported by the NCN grants 2022/01/1/ST2/00022 and 2023/49/B/ST2/03273, and by the AGH IDUB and the CERN COAS programmes.

### v2

@@ 2 The LHeC interaction region
It should be emphasised that the case described here refers to the truly concurrent operation of electron-hadron and hadron-hadron collisions. So, the aperture needed in the mini-beta quadrupoles of the protons is defined by both the colliding and the non-colliding proton beam. Therefore, that non-colliding proton beam still has to be kept in the storage ring and guided through the $ep$ interaction region untouched. However, in order to save free aperture for the optics of the colliding protons, relaxed beam optics has been developed which allows for minimum beam size and aperture need in the mini-beta quadrupoles: $\beta^{\ast}=24~{\rm m}$ has been found as optimum value for the non-colliding proton in this case.
[…]
Polarisation. Polarised electron beams can be rather naturally incorporated into an ERL type accelerator. Laser-driven state-of-the-art electron guns provide polarised beams up to 90% as previously suggested in the LHeC CDR [2]). Harmful effects like resonant depolarisation that play a major role in storage rings are by definition not existent, and the de-tuning of the spin precession due to the emitted synchrotron light is estimated to be on the order of a few percent. As a consequence the acceleration chain provided by the ERL is quasi-spin transparent and does not play a large role in the final – longitudinal – electron beam polarisation at the IP.
@@ 3 Physics case
In the phase-one LHeC, the centre-of-mass energy for electron-proton collisions $\sqrt{s_{ep}}=\sqrt{4E_{e}E_{p}}=0.75$ TeV, in place of 1.2 TeV for the 50 GeV electron beam at the final LHeC. This means that the kinematic reach on the $(x,Q^{2})$ plane will already be greatly extended with respect to the data taken at HERA at $\sqrt{s}=0.32$ TeV, as $Q^{2}_{\rm max}\simeq s_{ep}$ and $x_{\rm min}\propto 1/s_{ep}$. As discussed in the LHeC CDR [2], a relatively low integrated luminosity of $50~{\rm fb}^{-1}$ is sufficient to greatly improve the constraints on the parton density functions, particularly at low $x$. The longitudinal electron polarisations of both signs will be instrumental in decomposing the parton distributions, and in particular, in determining the total contribution of sea quarks and their possible charge asymmetry, as well as for measuring the quark density ratio $u/d$. One should stress that it also includes studies of the nuclear parton distributions. Therefore, already after two years of running the phase one LHeC it will be possible to provide valuable feedback to hadron-hadron experiments at the HL-LHC, which will result in significant improvements in a number of precision measurements [4]. Moreover, with entering into a new kinematic region of very low $x$ the novel QCD phenomena are expected to be observed – new parton dynamics and non-linear parton density effects.
Table 3: Inclusive Higgs, single top, and top pair generator level production cross sections for the phase-one and final LHeC, and for un-polarised and polarised electron beams. In the definition of the process, X corresponds to the dissociative state resulting from the proton interaction. All cross sections are quoted at leading order, with the Higgs boson mass of $125~{\rm GeV}$ and the top quark mass of $172~{\rm GeV}$.
@@
To illustrate the scientific potential of the research in electroweak and Higgs physics at the phase-one LHeC, in Tab. 3 we compare four generator-level cross-sections of processes of interest for un-polarised and longitudinally polarised electron beams to those for the final LHeC. All the values cited here are obtained using the version 3.5.8 of the MadGraph5_aMC@NLO [14] event generator, using the CTEQ6 [15] set of parton density functions evaluated at leading order. The Higgs boson case remains compelling at 0.75 TeV (especially for the charged current events), as well as the single-top-quark one, not only for mastering pioneering electron-proton measurement techniques but also for making measurements of some specific channels. For example, assuming the Run5 total integrated luminosity of $250~{\rm fb}^{-1}$ at 0.75 TeV, and a 80-80% electron longitudinal polarisation, we obtain the following uncertainties by scaling up the errors obtained for the 1.2 TeV LHeC [2] using the numbers of produced events: in the nomenclature of the $\kappa$ framework of an effective field theory (EFT) extension in the Higgs sector [16], $\delta\kappa_{W}\approx 2.5\%$ for the Higgs boson coupling to $W$ bosons, and $\delta{\cal R}e(f^{2}_{L})\approx 0.2$, for the anomalous $Wtb$ coupling in dimension-6 EFT extensions [17] – both errors comparable to the expected final uncertainties for the corresponding measurements at the HL-LHC using $3~{\rm ab}^{-1}$ of the $pp$ data.
@@
In Fig. 7 we show the differential cross-sections in rapidity for the Higgs boson and a single top quark produced at the phase-one LHeC as well as the final one – for the latter the cross-sections peak at about $Y=2$ requiring good coverage in the forward part of the hadron hemisphere. For the main decay products of these two exemplary processes of interest, a pseudo-rapidity coverage of up to 4 units, both in the tracking and calorimetry components of this hemisphere, would ensure a proper detection efficiency. The big advantage of using the highly polarised electron beam is clearly visible.
@@
Figure 7: Differential cross-sections in rapidity for the Higgs boson (left) and a single top quark (right), both produced through charged current, assuming the phase-one electron beam energy of 20 GeV, with and without electron polarisation, as well as for the final electron beam energy of 50 GeV.
@@
Experimental conditions at the LHeC will be particularly favourable for studies of the exclusive production via photon-photon fusion, the event pile-up will be negligible, highly improving the efficiency and purity of exclusive event selection, and at the same time some sources of overwhelming backgrounds, as due to the Drell-Yan process for example, will be absent. Moreover, no event suppression will occur due to hadronic re-scattering and the associated significant theoretical and experimental uncertainties will not intervene. Finally, if we consider the usage of data streaming (or the absence of triggering) in acquisition, the detection and reconstruction of a much wider range of final states will be allowed. As a result, despite the lower average $\gamma\gamma$ centre-of-mass energy than that at the HL-LHC, a significantly larger statistics of detected photon-photon interactions is expected already at the phase-one LHeC [18].
@@
For example, using the CepGen v1.2.2 [19] implementation of the LPAIR [20] matrix element-level event generator, we estimate a large cross-section for the fully exclusive two-photon production of tau-lepton pairs: about 40 pb at the phase-one LHeC, for pair invariant masses above 10 GeV. This will lead to very large event statistics within the acceptance of the central detector and as a consequence to excellent sensitivity to the $\tau$ anomalous magnetic moment $a_{\tau}$. With an integrated $ep$ luminosity of 250 fb-1, the expected sensitivity is an order of magnitude better than that achieved at the LHC [21]. As a result, for the first time the experimental uncertainties will be better than the higher-order corrections to the $\tau$ magnetic moment in the Standard Model. Similarly, large improvements will be achieved in constraining the electric dipole moment of tau-leptons.
Of course, there are many more searches for the BSM physics to be performed at the LHeC as, for example, unique searches for supersymmetric particles and studies of anomalous quartic gauge couplings [18]. Moreover, signatures such as displaced vertices or short tracks may emerge, which can be efficiently identified at the LHeC due to the low level of hadronic background and event pile-up with respect to the LHC, and as such offering a considerable window for discovery. Notable examples involve heavy neutrinos and axion-like particles, inspired by the QCD axion [2, 4].
@@
Last but not least, already at the phase-one LHeC, also one of the key questions in nuclear physics can be addressed: how the nucleon structure is modified when immersed in a nuclear medium. For example, by a precise measurement and complete unfolding of the distributions of all parton species in a single nucleus comparable to that in the proton, the LHeC will revolutionise our understanding of nuclear structure in crucial and unique kinematic regions [2, 4]. In addition, it will allow to clarify the dynamics for particle production in nuclear collisions in the region of relevance for the LHC. Finally, it should be stressed that the phase-one LHeC provides an important opportunity to perform $eA$ and $AA$ measurements using the same detector which provides the unique advantage for cross-calibration of performance and physics.
The overlap in the physics programmes of the LHeC and the Electron-Ion Collider (EIC) at Brookhaven is small, as the two projects will produce very important but also very complementary scientific outputs. The maximal centre-of-mass energy of electron-proton collisions at the EIC is 0.14 TeV therefore the electroweak and Higgs physics play there only marginal role, in stark contrast to the LHeC. Moreover, the significantly higher centre-of-mass energy at the LHeC corresponds to a much wider coverage on the $(x,Q^{2})$ plane, as explained above, allowing for unique studies of non-linear effects in partonic interactions and for significantly improved predictions of low-$x$ processes in the $pp$ collisions at the LHC. On the other hand, both beams can be polarised at the EIC and that will provide means for new insights into the origin of spin in hadrons, in particular. Last but not least, there is a great synergy between the two projects in the detector developments and of the associated electronics.
@@ 4 General detector requirements
In 2022, an extension of the ALICE experiment beyond Run4 at the HL-LHC was proposed in a letter of intent [22]. In that document, two configurations of magnetic field were considered for the new ALICE 3 detector at the IP2 – produced either by the cost effective 7.5 m long solenoid, or by the high-performance combination of a 2.5 m solenoid at the centre with two 2 m long dipoles on its each side. As discussed above, to allow for the optimal electron-hadron collisions, the dipole field is necessary for the separation of the electron beam, so the latter approach is desirable in this context.
The most recent ALICE 3 design includes a tracking detector with excellent performance that covers a wide range of rapidities of $|\eta|<4$ [23], which is well compatible with the needs for electron-hadron measurements at the LHeC with 20 GeV electrons. However, in contrast, the present design lacks a hadronic calorimeter (HCAL) while a performant one is obligatory for performing proper measurements of the DIS charged current events and reconstruction of the decays of top quark, Higgs, $W$ and $Z$ bosons. Moreover, an extension of the electromagnetic calorimeter is desirable in the electron forward direction to ensure a good reconstruction of DIS neutral current events; in addition, important studies of exclusive production at the LHeC would largely benefit from specialised near-beam detectors in very forward electron and proton directions.
Finally, it should be stressed that the ALICE 3 design already contains a full data-streaming mechanism (i.e. a ”trigger-less” data acquisition), which is also very desirable for measurements at the LHeC.
In summary, it seems that the proposed ALICE 3 detector requires relatively small modifications, apart from the addition of HCAL, to make it also appropriate for the phase-one LHeC experiment at the IP2 – in this way, allowing vastly increasing the scientific scope and impact of research performed using the Run5 data.
@@ 5 Conclusions
In this paper, we argue that a phased realisation of LHeC offers the best approach. Firstly, in this case the project financial as well as work load will be distributed over more years. Secondly, the experience acquired during the first stage of the project will help to make an optimal design of the final LHeC and possibly also allow its earlier high-performance start-up. For example, there are two technical items that will be studied and optimised during the phase one:
  • The performance of an ERL strongly depends on the understanding and control of the RF-phase during the acceleration and deceleration stages. Unlike for a synchrotron, where the synchronous phase is automatically adjusted, phase stability is of crucial importance here.
  • Experience at the SLAC and HERA colliders has shown that orbit fluctuations at the IP have to be damped to avoid emittance blow-up of the beams. The LHC itself is equipped with an orbit stabilisation system to avoid these effects, and experience will show whether additional systems will be needed for $ep$ collisions, e.g. multilayer orbit control as used in high-performance light sources or free-electron lasers.
The exploratory studies presented above demonstrated the high luminosity performance of such a phase-one LHeC. This in turn guarantees an excellent scientific output already during Run5: firstly, it will ensure important and timely scientific feedback to hadron-hadron experiments in HL-LHC as well as unique studies of new phenomena due to high density of partons; secondly, it will allow for making interesting contributions to the studies of electroweak and top physics, for example, and for performing original searches for the BSM physics; and finally, it will allow unique comparative studies of the $eA$ and $AA$ interactions, as both types of collisions are allowed by the proposed IR design.
The new detector at the P2 site can be designed by extension of the ALICE 3 proposal, which offers strong synergies and benefits in several areas.
@@ Acknowledgements
We thank T. M. P. von Witzleben for her contribution on the study of the beta-beat in the full HL-LHC lattice. L. Forthomme and K. Piotrzkowski appreciate the financial support of the Polish National Agency for Academic Exchange (NAWA) under grant number BPN/PPO/2021/1/00011. L. Forthomme has also been supported by the NCN grants 2022/01/1/ST2/00022 and 2023/49/B/ST2/03273, and by the AGH IDUB and the CERN COAS programmes.